Assume that all three retailers sell a popular shirt that retails for $32. To compare the impact of inventory costing method, we will also assume that all three retailers have the following inventory and sales data for the same period. To keep the calculations simple, a “unit” represents one million shirts.

· Beginning inventory: 3 units @ $10.00

· Purchases: 2 units @ $15.00

· Ending inventory: 1 unit

Questions

1. Calculate the number of units sold.

2. Calculate cost of goods sold, cost of ending inventory, and gross margin for Best Buy using the weighted-average inventory costing method.

3. Calculate cost of goods sold, cost of ending inventory, and gross margin for Amazon using the FIFO inventory costing method.

4. Calculate cost of goods sold, cost of ending inventory, and gross margin for Target using the LIFO inventory costing method.

5. If prices are decreasing, would FIFO or LIFO produce the highest income? Explain.

6. Would FIFO or LIFO produce the highest inventory cost on the balance sheet if prices are rising? Explain.
Answer #1
All numbers are in millions
1. Number of units sold: Opeing stock + purchases - closing stock = 3+2-1 = 4 units
2. Weighted average cost per unit: (3*10 + 2 *15) / (3+2) = 12
Cost of goods sold = No. of units sold X weighted averge cost per unit
Cost of goods sold = 4 unit x $ 12 = $ 48
Cost of ending inventory = Closing stock X Weighted average cost per unit
    Cost of ending inventory = 1 unit x $ 12 = $ 12
Gross margin = Sales revenue - cost of goods sold
    Gross margin = $ (32 * 4) - $ 48 = $ 80
3. Value of ending inventory under FIFO = Cost per unit of purchase x closing stock
    Value of ending inventory under FIFO = $ 15 X 1 = $ 15
Cost of goods sold = Total cost - cost of closing stock = (3*$ 10+2*$ 15) - $ 15 = $ 45
Gross margin = Sales revenue - cost of goods sold
    Gross margin = $ (32 * 4) - $ 45 = $ 83
4. Value of ending inventory under LIFO = Cost per unit in opening stock x closing stock
    Value of ending inventory under LIFO = $ 10 X 1 = $ 10
Cost of goods sold = Total cost - cost of closing stock = (3*$ 10+2*$ 15) - $ 10 = $ 50
Gross margin = Sales revenue - cost of goods sold
    Gross margin = $ (32 * 4) - $ 50 = $ 78
5. When prices are decreasing FIFO method will give higher income. This is because lower value cost will be charged to income statement and higher value stock will be lying in closing inventory.
6. If prices are rising, this means that last purchase will be at higher cost and under FIFO method last purchase will be part of closing inventory. So if prices are rising, FIFO method will give highest inventory cost.